# CyC3D: Fine-grained Controllable 3D Generation via Cycle Consistency Regularization

### v1

@@ 3.4 View Cycle Consistency Regularization
Rendering Regularization Loss: To regularize the possible failures in generation, the image reconstruction loss between the rendered images from generated 3D contents and the ground truth images on different views are used to supervise the generation model $f_{\text{gen}}$:
$\small L_{\text{render}}=\sum_{T_{l}}\|f_{\text{render}}(f_{\text{gen}}(C_{v},C_{i},T_{i}),T_{l})-I_{l}^{\text{gt}}\|_{2}^{2},$ (15)
@@
$\small L_{\text{clip}}=\sum_{T_{r}}(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\hat{I}_{r})))+$ (16)
@@
$\small L_{\text{view}}=L_{\text{render}}+\alpha\cdot L_{\text{clip}},$ (17)
where $\alpha$ is the set to 5.0 following [31] to balance the scale of different terms as default.
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN [6] 13.01 0.780 0.0526 15.07 0.772 0.0322 0.1277 0.1268 0.0311 0.0230 0.0497
GaussianDreamer [34] 12.41 0.766 0.0597 14.58 0.792 0.0362 0.0740 0.0947 0.0391 0.0225 0.0366
DreamGaussian [26] 9.94 0.686 0.1048 14.42 0.765 0.0380 0.0980 0.0916 0.0406 0.0216 0.0335
VolumeDiffusion [28] 13.25 0.818 0.0497 16.50 0.836 0.0231 0.1202 0.0953 0.0531 0.0190 0.0172
3DTopia [11] 9.63 0.704 0.1134 15.54 0.801 0.0287 0.1193 0.1064 0.0420 0.0268 0.0372
MVControl [15] 10.44 0.723 0.0952 14.51 0.746 0.0370 0.0555 0.0645 0.0160 0.0209 0.0315
ControLRM-T [31] 12.43 0.836 0.0607 17.17 0.836 0.0202 0.0567 0.0713 0.0599 0.0120 0.0359
ControLRM-D [31] 12.47 0.837 0.0600 17.23 0.837 0.0197 0.0383 0.0546 0.0650 0.0119 0.0357
Table 1: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on GSO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal. Our CyC3D-T/-D achieves significant controllability improvements in the table.
Methods Edge Sketch Depth Normal
PSNR SSIM MSE PSNR SSIM MSE M-MSE Z-MSE R-MSE NB-MSE DN-CON
GSGEN [6] 12.75 0.737 0.0614 14.25 0.744 0.0384 0.1004 0.0967 0.0463 0.0171 0.0479
GaussianDreamer [34] 11.92 0.736 0.0699 13.75 0.734 0.0439 0.0809 0.1061 0.0521 0.0188 0.0463
DreamGaussian [26] 9.38 0.683 0.1178 12.90 0.716 0.0523 0.1267 0.0762 0.0433 0.0188 0.0349
VolumeDiffusion [28] 12.58 0.734 0.0619 15.03 0.780 0.0321 0.1007 0.1198 0.0671 0.0188 0.0264
3DTopia [11] 8.48 0.633 0.1469 14.41 0.741 0.0372 0.1916 0.1844 0.0559 0.0346 0.0418
MVControl [15] 9.52 0.656 0.1161 13.34 0.686 0.0479 0.0454 0.0467 0.0181 0.0133 0.0350
ControLRM-T [31] 10.49 0.552 0.0909 15.77 0.777 0.0272 0.0514 0.0497 0.0700 0.0136 0.0572
ControLRM-D [31] 11.51 0.592 0.0725 15.80 0.778 0.0269 0.0493 0.0545 0.0661 0.0158 0.0548
Table 2: Quantitative comparison of controllability with state-of-the-art controllable 3D generation methods on ABO benchmark. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. The best results are highlighted in bold. We provide the results of four different conditions: edge, sketch, depth, and normal. Our CyC3D-T/-D achieves significant controllability improvements in the table.
@@ 4.1 Implementation Details
Datasets. Our training dataset is the training split of G-Objaverse (GOBJ) dataset [21], a sub-set of Objaverse [8]. Following the selection principle of [15], we utilize the test samples collected by [31] gathered from 2 different datasets for zero-shot evaluation: 80 samples from Google Scanned Objects dataset (GSO) [9], and 80 samples from Amazon Berkeley Objects dataset (ABO) [7].
[…]
@@
GSGEN [6] 344.61 0.740 0.289 366.47 0.669 0.259
GaussianDreamer [34] 278.70 0.810 0.300 225.38 0.787 0.277
DreamGaussian [26] 359.65 0.760 0.279 392.95 0.723 0.247
VolumeDiffusion [28] 299.61 0.715 0.259 350.46 0.679 0.242
3DTopia [11] 331.39 0.727 0.283 231.55 0.751 0.272
MVControl [15] 278.08 0.816 0.288 217.97 0.802 0.291
ControLRM-T [31] 260.75 0.846 0.289 202.14 0.827 0.282
ControLRM-D [31] 171.13 0.838 0.302 181.84 0.836 0.292
@@
Training Details. In analogy with [31], we initialize our network with the weights from pretrained OpenLRM-base [12]. For CyC3D-T, the conditional backbone is a transformer-based network; For CyC3D-D, the conditional backbone is a diffusion-based network. The training might cost 2 days for CyC3D-T and 3 days for CyC3D-D on 32 Nvidia V100-32G GPUs. More details are provided in the appendix.
Evaluation and Metrics. For the task of controllable 3D generation task, we consider 4 different conditions: edge, sketch, depth, and normal. For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated 3D contents following [31]. For edge/sketch condition, we utilize PSNR, SSIM, and MSE as evaluation metrics; For depth/normal condition, we utilize MSE to measure the similarity. In analogy with [14], the large foundation models in monocular depth estimation are used to capture the depth priors for model-based evaluation. If Midas [22] is used, the metric is noted as M-MSE. If ZoeDepth [3] is used, the metric is noted as Z-MSE. Following [31], the depth consistency in 3D space compares the difference between rendered depth map and the input condition depth map, noted as R-MSE. In analogy with [14], the metric of NB-MSE uses the pre-trained model (Normal-BAE [1]) in surface normal estimation to assess the normal consistency. Furthermore, we also evaluate the depth-normal consistency between the rendered depth map and the input normal maps following [31], noted as DN-CON. Please refer to more detailed introduction of these evaluation metrics in the appendix.
@@ 4.2 Experimental Results
ControlNet+OpenLRM-S [12] 0.1069 0.0459 0.1333 0.0255 268.7 0.807 0.292
ControlNet+OpenLRM-B [12] 0.0897 0.0452 0.1312 0.0206 258.04 0.822 0.293
ControlNet+OpenLRM-L [12] 0.0792 0.0447 0.1383 0.0226 256.43 0.824 0.294
ControlNet+TGS [36] 0.0523 0.0417 0.1403 0.0208 258.86 0.841 0.295
ControlNet+TripoSR [29] 0.0978 0.0421 0.1188 0.0241 261.76 0.835 0.289
MVControlNet+LGM [15] 0.0952 0.0370 0.0160 0.0315 278.08 0.816 0.288
@@
Qualitative Results. We provide qualitative results of our CyC3D in Fig. 4. The backbone is inherited from ControLRM [31], and the diffusion model is replaced with a pre-trained ControlNet. As the figure shows, our method not only performs well in generating hight quality 3D contents but also in preserving details in conditional controls. Furthermore, we also compare the qualitative results of our CyC3D with other state-of-the-art methods including ControLRM [31] and MVControl [15] in Fig. 5.
Controllability Evaluation. The comparisons of controllability with state-of-the-art controllable 3D generation methods on GSO/ABO benchmarks are presented in Tab. 1/2. Following [31], we report the controllability metrics discussed in Sec. 4.1 on all conditional controls. From the tables, we can find that our CyC3D can achieve state-of-the-art performance in the quantitative comparison with other baselines. For example, the R-MSE of our CyC3D-D is 0.0060, which is much lower than the baselines of ControLRM-D (0.0650 R-MSE). Furthermore, we also provide qualitative comparison of controllability in Fig. 6.
Generation Evaluation. To investigate whether improving controllability compromises generation quality, we provide the quantitative results of Fréchet Inception Distance (FID) [10] and CLIP-score [18] following [15] in Tab. 3. We report CLIP-I metric to quantify image similarity between reference and rendered images, and CLIP-T metric to evaluate text-image alignment using textual annotations. In rigorous zero-shot generalization experiments of cross-domain transitions towards GSO and ABO, our proposed CyC3D demonstrates statistically significant improvements over existing methods. Furthermore, from the qualitative results in Fig. 5, out CyC3D can generate better results with higher quality and consistency than existing methods.
Comparison with ControlNet-based Baselines. For a fair comparison, we also provide experimental results compared with ControlNet combined with image-to-3D reconstruction models. The images produced by ControlNet are processed through image-to-3D reconstruction models to synthesize 3D assets, with foreground segmentation implemented using Rembg for background removal. The quantitative evaluation results on the GSO benchmark are presented in Tab. 4, incorporating state-of-the-art image-to-3D systems including OpenLRM [12], TGS [36], and TripoSR [29] as comparative baselines.
@@ 4.3 Ablation Study
Effect of Condition Cycle Consistency. To evaluate the effectiveness of 2D conditional control feedback (e.g. Eq. 10 for edge/sketch), we provide ablation experiments in Tab. 5. Furthermore, we also provide ablation results of 3D conditional control feedback (Eq. 12 for depth, Eq. 14 for normal) in Tab. 6. As the tables show, each term of condition cycle consistency can effectively imporve the controllability upon different conditional controls.
Effect of View Cycle Consistency. To evaluate the effectiveness of view cycle consistency feedback (e.g. Eq. 17), we provide qualitative ablation results in Fig. 7. As the figure shows, without view cycle consistency regularization, the generated 3D contents might be overfitting to the input viewpoint. Though the rendered results in the input view is reasonable, the lack of regularization on the remaining views might lead to a low-quality 3D shape.

### v2

@@ 3.4 View Cycle Consistency Regularization
Rendering Regularization Loss. To regularize the possible failures in generation, the image reconstruction loss between the rendered images from generated 3D contents and the ground truth images on different views is used to supervise the generation model $f_{\text{gen}}$:
$L_{\text{render}}=\sum_{T_{l}}\|f_{\text{render}}(f_{\text{gen}}(C_{v},C_{i},T_{i}),T_{l})-I_{l}^{\text{gt}}\|_{2}^{2},$ (15)
@@
$L_{\text{clip}}=\sum_{T_{r}}(1-\cos(f_{\text{clip-t}}(C_{t}),f_{\text{clip-i}}(\hat{I}_{r})))+$ (16)
@@
$L_{\text{view}}=L_{\text{render}}+\alpha\cdot L_{\text{clip}},$ (17)
where $\alpha$ is the set to 5.0 following (Xu et al. 2024a) to balance the scale of different terms as default.
@@ 4.1 Implementation Details
Datasets. Our training dataset is the training split of the G-Objaverse (GOBJ) dataset (Qiu et al. 2024), a sub-set of Objaverse (Deitke et al. 2023). Following the selection principle of (Li et al. 2024), we utilize the test samples collected by (Xu et al. 2024a) gathered from 2 different datasets for zero-shot evaluation: 80 samples from Google Scanned Objects dataset (GSO) (Downs et al. 2022), and 80 samples from Amazon Berkeley Objects dataset (ABO) (Collins et al. 2022).
Methods $\rightarrowGSO GOBJ\rightarrowABO$
@@
GSGEN (Chen et al. 2024) 344.61 0.740 0.289 366.47 0.669 0.259
GaussianDreamer (Yi et al. 2024) 278.70 0.810 0.300 225.38 0.787 0.277
DreamGaussian (Tang et al. 2023a) 359.65 0.760 0.279 392.95 0.723 0.247
VolumeDiffusion (Tang et al. 2023b) 299.61 0.715 0.259 350.46 0.679 0.242
3DTopia (Hong et al. 2024) 331.39 0.727 0.283 231.55 0.751 0.272
MVControl (Li et al. 2024) 278.08 0.816 0.288 217.97 0.802 0.291
ControLRM-T (Xu et al. 2024a) 260.75 0.846 0.289 202.14 0.827 0.282
ControLRM-D (Xu et al. 2024a) 171.13 0.838 0.302 181.84 0.836 0.292
@@
Training Details. In analogy with (Xu et al. 2024a), we initialize our network with the weights from pretrained OpenLRM-base (Hong et al. 2023). For CyC3D-T, the conditional backbone is a transformer-based network; For CyC3D-D, the conditional backbone is a diffusion-based network. The training might cost 2 days for CyC3D-T and 3 days for CyC3D-D on 32 Nvidia V100-32G GPUs. More details are provided in the appendix.
Evaluation and Metrics. For the task of controllable 3D generation task, we consider 4 different conditions: edge, sketch, depth, and normal. For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated 3D contents following (Xu et al. 2024a). For edge/sketch condition, we utilize PSNR, SSIM, and MSE as evaluation metrics; For depth/normal condition, we utilize MSE to measure the similarity. In analogy with (Li et al. 2025), the large foundation models in monocular depth estimation are used to capture the depth priors for model-based evaluation. If Midas (Ranftl et al. 2020) is used, the metric is noted as M-MSE. If ZoeDepth (Bhat et al. 2023) is used, the metric is noted as Z-MSE. Following (Xu et al. 2024a), the depth consistency in 3D space compares the difference between the rendered depth map and the input condition depth map, noted as R-MSE. In analogy with (Li et al. 2025), the metric of NB-MSE uses the pre-trained model (Normal-BAE (Bae, Budvytis, and Cipolla 2021)) in surface normal estimation to assess the normal consistency. Furthermore, we also evaluate the depth-normal consistency between the rendered depth map and the input normal maps following (Xu et al. 2024a), noted as DN-CON. Please refer to more details in the appendix.
@@ 4.2 Experimental Results
ControlNet+OpenLRM-S (Hong et al. 2023) 0.1069 0.0459 0.1333 0.0255 268.7 0.807 0.292
ControlNet+OpenLRM-B (Hong et al. 2023) 0.0897 0.0452 0.1312 0.0206 258.04 0.822 0.293
ControlNet+OpenLRM-L (Hong et al. 2023) 0.0792 0.0447 0.1383 0.0226 256.43 0.824 0.294
ControlNet+TGS (Zou et al. 2024) 0.0523 0.0417 0.1403 0.0208 258.86 0.841 0.295
ControlNet+TripoSR (Tochilkin et al. 2024) 0.0978 0.0421 0.1188 0.0241 261.76 0.835 0.289
MVControlNet+LGM (Li et al. 2024) 0.0952 0.0370 0.0160 0.0315 278.08 0.816 0.288
@@
Qualitative Results. We provide qualitative results of our CyC3D in Fig. 10. The backbone is inherited from ControLRM (Xu et al. 2024a), and the diffusion model is replaced with a pre-trained ControlNet. As the figure shows, our method not only performs well in generating high-quality 3D content but also in preserving details in conditional controls. Furthermore, we also compare the qualitative results of our CyC3D with other state-of-the-art methods, including ControLRM (Xu et al. 2024a) and MVControl (Li et al. 2024) in Fig. 4.
Controllability Evaluation. The comparisons of controllability with state-of-the-art controllable 3D generation methods on GSO/ABO benchmarks are presented in Tab. 1/2. Following (Xu et al. 2024a), we report the controllability metrics discussed in Sec. 4.1 on all conditional controls. From the tables, we can find that our CyC3D can achieve state-of-the-art performance in the quantitative comparison with other baselines. For example, the R-MSE of our CyC3D-D is 0.0060, which is much lower than the baselines of ControLRM-D (0.0650 R-MSE). Furthermore, we also provide a qualitative comparison of controllability in Fig. 5.
Generation Evaluation. To investigate whether improving controllability compromises generation quality, we provide the quantitative results of Fréchet Inception Distance (FID) (Heusel et al. 2017) and CLIP-score (Mohammad Khalid et al. 2022) following (Li et al. 2024) in Tab. 3. Furthermore, the qualitative results are provided in Fig. 4. We report the CLIP-I metric to quantify image similarity between reference and rendered images, and the CLIP-T metric to evaluate text-image alignment using textual annotations. In rigorous zero-shot generalization experiments of cross-domain transitions towards GSO and ABO, our proposed CyC3D demonstrates statistically significant improvements over existing methods.
Comparison with ControlNet-based Baselines. For a fair comparison, we also provide experimental results compared with ControlNet combined with image-to-3D reconstruction models. The images produced by ControlNet are processed through image-to-3D reconstruction models to synthesize 3D assets, with foreground segmentation implemented using Rembg for background removal. The quantitative evaluation results on the GSO benchmark are presented in Tab. 4, incorporating state-of-the-art image-to-3D systems including OpenLRM (Hong et al. 2023), TGS (Zou et al. 2024), and TripoSR (Tochilkin et al. 2024) as comparative baselines.
@@ 4.3 Ablation Study
Effect of Condition Cycle Consistency. To evaluate the effectiveness of 2D conditional control feedback (e.g., Eq. 10 for edge/sketch), we provide ablation experiments in Tab. 5. Furthermore, we also provide ablation results of 3D conditional control feedback (Eq. 12 for depth, Eq. 14 for normal) in Tab. 6. As the tables show, each term of condition cycle consistency can effectively improve the controllability upon different conditional controls.
Effect of View Cycle Consistency. To evaluate the effectiveness of view cycle consistency feedback (e.g., Eq. 17), we provide qualitative ablation results in Fig. 6. As the figure shows, without view cycle consistency regularization, the generated 3D contents might be overfitting to the input viewpoint. Though the rendered results in the input view are reasonable, the lack of regularization on the remaining views might lead to a low-quality 3D shape.